OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find f(2) + g(3).

OpenStudy (anonymous):

f(2)=2(2)+1=5 g(3)=3(3)+1=10 5+10=15

OpenStudy (anonymous):

thank you so much